The prelim2e package^{∗}

Martin Schröder

Crüsemannallee 3 28213 Bremen Germany martin@oneiros.de

                               PGP-Key:  2048 bit / KeyID 292814E5
                            7E86 6EC8 97FA 2995  82C3 FEA5 2719 090E

2001/02/17

Abstract

This package allows the marking of (preliminary) versions of a document on the output.

Contents

1 Introduction 1

2 Usage 1

3 Options 2 4 Required packages 2

5 Putting more things at the bottom 3

6 Acknowledgements 3

1 Introduction

This package allows the marking of (preliminary) versions of a document. This is done using the command \PrelimText, whose expansion is added below the footer of every page of a document (look at the bottom of this page for an example).

2 Usage

Simply using this package via \usepackage{prelim2e} produces a text in the form of “Preliminary version – February 28, 2003”.

\PrelimText The text is produced by the commands \PrelimText and \PrelimWords, which \PrelimWords can be changed via \renewcommand or by using options at the \usepackage com-

mand (see section 3).

The footer for this documentation was produced by changing \PrelimWords to

1

∗The version number of this file is v1.23, last revised 2001/02/17.

prelim2e package – Version v1.23 – Documentation L^{A}TEXed – February 28, 2003 – 11:51

     \renewcommand{\PrelimWords}{%
        \textsf{prelim2e} package --
        Version \fileversion{} --
        Documentation \LaTeX{}ed%
        }

3 Options

The package has the following options: danish This sets the text to “Foreløbig version”. It does not use the babel package.

New feature 2001/02/17

draft If this option is used a text appears below the normal pagebody. It is the

default. final If this option is used prelim2e produces no text. english This sets the text to “Preliminary version”. It is the default. french This sets the text to “Version préliminaire”. It does not use the french or

New feature 1997/05/12

babel package. german This sets the text to “Vorläufige Version”. It does not use the german or

babel package. italian This sets the text to “Versione preliminare”. It does not use the babel

New feature 2001/02/17

package.

New description time This adds the time at the beginning of the current L^{A}TEX to the text. The 1997/05/12 time is produced using the \thistime command, which may be provided

by the scrtime package[2]. If you don’t use the scrtime package (e. g. by specifying the scrtime option), you must provide it yourself or by some other package.

scrtime This loads the scrtime package[2] (part of the KOMA-Script bundle), which

New feature 1997/05/12

provides the \thistime command used by the time option. The scrtime option implies the time option.

All other options are passed to the scrtime package if the scrtime option is selected.

4 Required packages

The package requires the following packages: everyshi[3] It is used to implement the setting of the text below the normal page-

body.

New description scrtime[2] It is used to typeset the current time and is needed if the scrtime option 1997/05/12 is selected.

5 Putting more things at the bottom

Another package you might like to use with prelim2e is vrsion[4]. This allows the definition and maintenance of a version number like 3.14159 within L^{A}TEX, which can be put at the bottom of every page using prelime2e.

6 Acknowledgements

The idea of this package is based on prelim.sty[1] for L^{A}TEX2.09 by Robert Tolksdorf (tolk@cs.tu-berlin.de). It provides nearly the same functionality as prelim2e, but in a very dirty way: it uses a modified output routine and does not work with L^{A}TEX 2ε.

The time functionality now uses the scrtime package by Markus Kohm

(markus.kohm@gmx.de).
   Rowland Bartlett (R.BARTLETT@liverpool-john-moores.ac.uk) brought forth

the idea of separating the time option and the scrtime package; Daniel Courjon (dcourjon@utinam.univ-fcomte.fr) provided the text for the french option, Arne Jørgensen (arne.jorgensen@tug.dk) provided the text for the danish option, and Davide Giovanni Maria Salvetti (salve@debian.org) and Riccardo Murri (murri@phc.unipi.it) both provided the text for the italian option,

As usual Rebecca Stiels improved the quality of this documentation.

References

[1] Mats Dahlgren. vrsion – a L^{A}TEX Macro for version Numbering of Files. CTAN:

   tex-archive/macros/latex/contrib/supported/vrsion/vrsion.dtx.

L^{A}TEX 2ε package.

[2] Markus Kohm. The KOMA-timedate-bundle. CTAN:tex-archive/macros/

   latex/contrib/supported/koma-script/scrtime.dtx. L^{A}TEX 2ε package.

[3] Martin Schröder. The everyshi package.

   CTAN:tex-archive/macros/latex/contrib/supported/ms/everyshi.dtx.

L^{A}TEX 2ε package.

[4] Robert Tolksdorf. Kennzeichnung von Vorversionen eines Dokuments.

   CTAN:tex-archive/macros/latex209/contrib/prelim/. L^{A}TEX2.09

package.